OpenStudy (anonymous):

help pleaasssssee 1/2q = 8

OpenStudy (anonymous):

is this \[\frac{q}{2} = 8\] or \[\frac{1}{2q} = 8\]

OpenStudy (anonymous):

either way, in order to get rid of the denominator on the left hand side you need to cross multiply. For example \[\frac{a}{b}=\frac{c}{d}\]is equivalent to\[a*d=c*b\]